# Interpretability of linear regression models of glassy dynamics

### v1

@@ IV Least square regression
is the mean square error (MSE) and $\mathcal{L}^{\rm reg}$ is a regularization term. In this section, we will show results for two standard linear models [70], namely ordinary least square regression (Sec. IV.1) and ridge regression (Sec. IV.2), and discuss their limitations.
@@ IV.1.1 Definition
In ordinary least square (OLS) regression, the weight vector $\hat{\bf w}$ is determined by minimizing the MSE, i.e., $\mathcal{L}=\mathcal{L}^{\rm MSE}$. Setting $\nabla_{\hat{\bf w}}\mathcal{L}^{\rm MSE}(\hat{\bf w})={\bf 0}$ and remembering that all the features are normalized, the solution for the OLS regression (see the Appendix B for the derivation) is
@@ IV.1.2 Oscillatory behavior of the weights
Interestingly, however, the Pearson coefficient between the ground truth $Y$ and the prediction $\hat{Y}$ of OLS regression is very good, $R[{Y},{\hat{Y}}]\approx 0.89$, on either the train or the test datasets. These results represent a paradigmatic machine learning case in which the prediction accuracy is very good, but the interpretability is poor. We confirmed that the other descriptors, namely the SLO and JBB descriptors, also demonstrate this pathological behavior.
@@ IV.2.1 Definition
includes a regularization term that penalizes solutions with large weights. The parameter $\alpha$ controls the magnitude of the regularization. Setting $\nabla_{\hat{\bf w}}\mathcal{L}^{\rm Ridge}(\hat{\bf w})={\bf 0}$ yields the estimated weights (see the Appendix B for the derivation)
@@
Figure 4: Normalized weights as a function of $\alpha$ corresponding to (a) the radial features $G^{S}(k)$ for $0\leq k\leq 47$ and (b) the angular features $\Psi^{SS}(k)$ for $0\leq k\leq 21$. The width of the shaded areas corresponds to the standard deviation estimated over independent random training sets. The color code indicates the feature index $k$.
@@ IV.2.3 Prediction accuracy
Figure 5 displays the performance metrics for the BP descriptor. We show results obtained on both the test set or the train set, to verify the lack of overfitting. Both metrics indicate very good performance accuracy and display a flat maximum stretching over several orders of magnitudes in $\alpha$. The performance starts to degrade appreciably only when $\alpha>0.1$, as can be seen from the similar drops of $R$ and $R_{2}$. We found quantitatively similar results for the physically motivated descriptors introduced in Sec. III.4 (not shown). The comparison between the chaotic traces of Fig. 4 and the features performance metrics of Fig. 5 is striking. Since very different solutions (see Fig. 4) yield nearly identical prediction accuracies (see Fig. 5), it is clear that prediction alone cannot be taken as a criterion to choose $\alpha$. This delicate aspect was overlooked in previous studies [31, 46], where the regularization parameter $\alpha$ was chosen to maximize the correlation. These issues are a manifestation of multicollinearity in the dataset and leave us with the question how to define an optimal model.
@@ V.1.1 Definition
We start by quantifying the degree of multicollinearity using the condition number $\kappa(\mathrm{C})$ of the correlation matrix $\mathrm{C}$. A brief introduction to the condition number and its meaning is provided in the Appendix C. In a nutshell, $\kappa(\mathrm{C})$ provides an upper bound for the relative error in the solution of a linear problem when small statistical perturbations are present in the data. When $\kappa(\mathrm{C})$ is small, $\mathrm{C}$ is well-conditioned and the solution is stable. When $\kappa(\mathrm{C})$ is large, the matrix $\mathrm{C}$ is ill-conditioned and the error in the solution may become significant.
@@ V.1.2 Two-features model
where $r=R[{X^{(1)}},{X^{(2)}}]$ is the Pearson coefficient between the two features. When $r=0$, the two features are orthogonal, and $\mathrm{C}$ reduces to the identity matrix. When $r\to 1$ ($r\to-1$), the two features are perfectly correlated (anti-correlated), and hence ${\bf X}^{(1)}$ and ${\bf X}^{(2)}$ are linearly dependent.
@@ V.1.3 Origin of the oscillatory behavior of weights
We now seek to explain the oscillatory behavior of the weights observed in Fig. 3. This behavior arises from the ill-conditioning of the correlation matrix, due to the vanishing of its smallest eigenvalue $\lambda_{\rm min}$.
@@
V.2 Least square regression revisited
@@
We can now build on the multicollinearity analysis to identify a range of optimal solutions within Ridge regression. To see how the regularization term in Eq. (23) mitigates the effect of collinearity, we compute the condition number in the Ridge setting. We get
@@
which avoids the divergence of $\kappa(\mathrm{C}+\alpha\mathrm{I})$ when multicollinearity is severe ($\lambda_{\rm min}\to 0$). In Fig. 6(a), we show $\kappa(\mathrm{C}+\alpha\mathrm{I})$ as a function of $\alpha$. In the range $\lambda_{\textrm{min}}\ll\alpha\ll\lambda_{\textrm{max}}$, the condition numbers decreases like $\lambda_{\textrm{max}}\alpha^{-1}$. To reduce the effects of multicollinearity to an acceptable degree, one should lower the condition numbers down to at least $10^{3}$ or less [48]. This occurs around $\alpha=0.1$. The regularization becomes meaningless for $\alpha\gtrsim\lambda_{\rm max}$, since it entirely suppresses correlations in $\mathrm{C}$. The crossover of $\kappa$ towards 1 occurs when $\alpha$ becomes of the order of $\lambda_{\textrm{max}}\approx 100$, as confirmed by the Fig. 6(a). Since values of $\alpha$ in the range $0.1\lesssim\alpha\lesssim 100$ are acceptable but the performance accuracy decreases systematically in this range, see Fig. 5, we tentatively choose $\alpha=0.1$ as an optimal value. Finally, one can also see how the oscillatory behavior originating from multicollinearity is suppressed in Ridge regression. This point can already be grasped by comparing Fig. 3(a) and Fig. 3(b), but is best understood in the principal component basis, see the Appendix D.3.
@@ VI Towards interpretable linear regression models
The analysis presented in Sec. V shows that Ridge regression is an effective method to cope with the effects of multicollinearity in structure-dynamics datasets. However, the solutions of this linear model are still too high-dimensional to be physically interpretable. To achieve a simple physical picture, we must substantially reduce the dimensionality of the problem at hand. In this section, we consider a generalization of Ridge regression, called elastic net [72], that provides means to select the most relevant features, while reducing multicollinearity (Sec. VI.1). As an alternative approach, we use principal component regression, which builds upon a simple linear transformation of original features (Sec. VI.2).
VI.1 Elastic net regression
@@ VI.1.1 Definition
Figure 7: Elastic net regression of the dynamic propensity using the BP descriptor. Main panel: Pearson coefficient $R[Y,\hat{Y}$] for the optimal models (large circles) as a function of the number of selected features $P$. The values of $\alpha$ and $\beta$ corresponding to optimal models with $P\leq 10$ are indicated in the panel. The small circles indicate results for values $(\alpha,\beta)$ corresponding to sub-optimal models. Inset: number of selected features $P$ as a function of $\beta$ for selected values of $\alpha$.
VI.1.2 Selection of parsimonious models
To get a feeling of the role of $\beta$, we show in the inset of Fig. 7 the dependence of the number of selected features $P$ on $\beta$, for a few fixed values of $\alpha$. The analysis is performed as usual on the test set. As expected, the number of selected features decreases as the strength of the L1 regularization increases, reaching its minimum for $\beta=1$. A similar effect is observed with increasing $\alpha$.
Of course, the prediction performance of the model will vary with $\alpha$ and $\beta$. To define a set of optimal models that provide the best prediction accuracy on the test set, while minimizing the number of selected features, we proceed as follows. For each fixed value of $\alpha$, we perform elastic net regression starting from $\beta=1$ (Lasso regression). If the algorithm converges to a solution, we have found the best model for that value of $\alpha$. Otherwise, we reduce $\beta$ until a solution is found, and extract the corresponding value of $P$. We then compute the Pearson coefficient $R$ with the dynamic propensity for each optimal model at a given $\alpha$.
The results of this procedure are shown in the main panel of Fig. 7 as large black symbols. We also include the results of a more extensive sampling of $(\alpha,\beta)$ pairs, shown as small dots. We see that the optimal models identified by our procedure always provide the best performance for a given $P$. The envelope of $R(P)$ defines a set of optimal models, in the sense that they provide the best performance accuracy for a given size $P$ of the feature space. The typical values of the Pearson coefficients for the most parsimonious models range from 0.54 for $P=1$, to 0.62 for $P=5$, to 0.74 for $P=10$. These values can be compared with the maximum value of 0.88 achieved when retaining the full descriptor, see also Sec. VII. Within our strategy, we found that pure Lasso regression ($\beta=1$) yields the optimal model, except in the case $M=1$.
@@ VI.1.2 Selection of parsimonious models
Model Pearson coefficient 0.53 0.53 0.53 0.51 0.47 0.39 0.35 0.31 0.28 -0.20 -0.28 -0.42
$P=10$ 0.74 $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
$P=8$ 0.69 $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
$P=5$ 0.62 $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
$P=3$ 0.60 $\bullet$ $\bullet$ $\bullet$
$P=2$ 0.58 $\bullet$ $\bullet$
$P=1$ 0.54 $\bullet$
@@
Table 2: Optimal elastic net regression models using the BP descriptor. The bullets indicate the features selected by each model for a given number of selected features $P$. The numbers in italic below each feature are the Pearson coefficients between that feature and the dynamic propensity. Only models retaining up to $P=10$ are shown.
@@
In Table 2, we summarize the features selected by parsimonious models ($P\leq 10$) along the envelope of Fig. 7. We include the corresponding values of the Pearson coefficients with the dynamic propensity, both for the regression models and for the individual features. We see that the low-dimensional models of elastic net regression (Lasso regression for $P\geq 2$) successfully pinpoint the features that are most correlated to the dynamic propensity, found in either angular or radial sectors of the BP descriptor. We notice, however, that these low-dimensional models are still somewhat redundant. For instance, the Pearson coefficient between the two features selected by the $P=2$ model is appreciable, $R=0.62$. Even worse, the model with $P=5$ contains two nearly identical angular features (163 and 164), having mutual correlation $R=0.99$. This can be explained by the fact that for these models $\beta=1$, hence, in practice, there is no suppression of multicollinearity. It would be possible to address this issue by using a minimum-redudancy-maximum-relevance selection scheme [74], which includes iteratively features that are highly correlated with the target, while penalizing those strongly correlated with the ones already taken. We leave this approach for a future investigation.
Although our numerical results indicate that the method is not able to fully remove multicollinearity, elastic net (Lasso) regression does overall a good job at selecting parsimonious models, achieving a reasonable correlation with the dynamic propensity ($R\approx 0.6$) using just a couple of features. The best performing low-dimensional model ($R\approx 0.74$), mixing both correlated and anti-correlated features, is obtained with $P=10$.
@@ VI.2.1 Definition
$\frac{1}{N_{\mathcal{S}}}{\mathrm{X}^{\prime}}^{T}\mathrm{X}^{\prime}=\mathrm{U}^{T}\mathrm{C}\mathrm{U}=\mathrm{\Lambda},$
@@
Figure 8: Principal component analysis for the BP descriptor: (a) eigenvalue $\lambda^{(f)}$ of each PC; (b) Pearson coefficient $R[\tilde{X}^{(f)},Y]$ between the dynamic propensity and each PC; (c) absolute value of the Pearson coefficient $|R[\tilde{X}^{(f)},Y]|$ between the dynamic propensity and each PC as a function of the eigenvalue $\lambda^{(f)}$ of the PC; (d) same as (c) but showing the eigenvalues in log scale. In panels (c) and (d), the top 5 PCs are marked by arrows.
@@ VI.2.2 Interpretation of the principal components
In Table 3 we report the Pearson coefficients between the projections on the first few PCs and the physically motivated features defined in Sec. III.4. Fluctuations along PC1 are moderately correlated to concomitant fluctuations of the local potential energy $\overline{u}$ and the bond orientational order parameter $\overline{\Psi}_{6}$. This structural mode gathers the bulk of the variance of the normalized dataset, but it is uncorrelated to the dynamic propensity. The interpretation of the second PC is sharp: structural fluctuations along PC2 are strongly correlated to the local number density $\overline{\rho}$ ($|R|=0.93$). Note that, as anticipated in Sec. III.6, in our model the fluctuations of the local packing fraction $\overline{\varphi}$ are strongly anti-correlated with those of $\overline{\rho}$. The remaining PCs do not seem to have clear connections with physically motivated features. Interestingly, none of the structural modes identified by PCA captures the structural fluctuations associated to $\overline{\Theta}$, which is most strongly correlated with the dynamic propensity. This suggests a limitation of the BP descriptor in accounting for subtle structural correlations related to local packing efficiency.
Pearson coefficients
$Y$ $\overline{\Theta}$ $\overline{u}$ $\overline{\rho}$ $\overline{\Psi}_{6}$ $\overline{\varphi}$ $\overline{z}$
$\tilde{X}^{(1)}$ +0.07 +0.06 +0.53 +0.07 +0.52 -0.01 -0.03
$\tilde{X}^{(2)}$ -0.43 -0.14 +0.13 -0.93 +0.30 +0.86 +0.14
$\tilde{X}^{(3)}$ -0.34 -0.33 -0.44 +0.00 -0.28 -0.04 +0.07
$\tilde{X}^{(4)}$ +0.31 +0.23 +0.31 +0.06 +0.09 -0.02 -0.07
$\tilde{X}^{(5)}$ -0.21 -0.23 -0.17 +0.04 -0.07 -0.03 +0.05
Table 3: Pearson coefficients $R$ between the first five PCs and physically motivated features coarse-grained over a length $\ell=1.5$. Values of $R$ larger than 0.5 in absolute value are shown in bold.

### v2

@@ IV Least square regression
is the mean square error (MSE) and $\mathcal{L}^{\rm reg}$ is a regularization term. In this section, we will show results for two standard linear models [70], namely ordinary least squares regression (Sec. IV.1) and Ridge regression (Sec. IV.2), and discuss their limitations.
@@ IV.1.1 Definition
In ordinary least squares (OLS) regression, the weight vector $\hat{\bf w}$ is determined by minimizing the MSE, i.e., $\mathcal{L}=\mathcal{L}^{\rm MSE}$. Setting $\nabla_{\hat{\bf w}}\mathcal{L}^{\rm MSE}(\hat{\bf w})={\bf 0}$ and remembering that all the features are normalized, the solution for the OLS regression (see the Appendix C for the derivation) is
@@ IV.1.2 Oscillatory behavior of the weights
Interestingly, however, the Pearson coefficient between the ground truth $Y$ and the prediction $\hat{Y}$ of OLS regression is very good, $R[{Y},{\hat{Y}}]\approx 0.87$, on either the train or the test datasets. These results represent a paradigmatic machine learning case in which the prediction accuracy is very good, but the interpretability is poor. We confirmed that the other descriptors, namely the SLO and JBB descriptors, also demonstrate this pathological behavior.
@@ IV.2.1 Definition
includes a regularization term that penalizes solutions with large weights. The parameter $\alpha$ controls the magnitude of the regularization. Setting $\nabla_{\hat{\bf w}}\mathcal{L}^{\rm Ridge}(\hat{\bf w})={\bf 0}$ yields the estimated weights (see the Appendix C for the derivation)
@@
Figure 4: Normalized weights as a function of the Ridge regularization parameter $\alpha$ corresponding to (a) the radial features $G^{S}(k)$ for $0\leq k\leq 47$ and (b) the angular features $\Psi^{SS}(k)$ for $0\leq k\leq 21$. The width of the shaded areas corresponds to the standard deviation estimated over independent random training sets. The color code indicates the feature index $k$.
@@ IV.2.3 Prediction accuracy
Figure 5 displays the performance metrics for the BP descriptor. We show results obtained on both the test set or the train set, to verify the lack of overfitting. Both metrics indicate very good performance accuracy and display a flat maximum stretching over several orders of magnitudes in $\alpha$. The performance starts to degrade appreciably only when $\alpha>0.1$, as can be seen from the similar drops of $R$ and $R_{2}$. We found quantitatively similar results for the physically motivated descriptors introduced in Sec. III.4 (not shown). The comparison between the chaotic traces of Fig. 4 and the features performance metrics of Fig. 5 is striking. Since very different solutions (see Fig. 4) yield nearly identical prediction accuracies (see Fig. 5), it is clear that prediction alone cannot be taken as a criterion to choose $\alpha$. This delicate aspect was overlooked in previous studies [31, 46, 63], where the regularization parameter $\alpha$ was chosen to maximize the correlation. These issues are a manifestation of multicollinearity in the dataset and leave us with the question how to define an optimal model.
@@ V.1.1 Definition
We start by quantifying the degree of multicollinearity using the condition number $\kappa(\mathrm{C})$ of the correlation matrix $\mathrm{C}$. A brief introduction to the condition number and its meaning is provided in the Appendix D. In a nutshell, $\kappa(\mathrm{C})$ provides an upper bound for the relative error in the solution of a linear problem when small statistical perturbations are present in the data. When $\kappa(\mathrm{C})$ is small, $\mathrm{C}$ is well-conditioned and the solution is stable. When $\kappa(\mathrm{C})$ is large, the matrix $\mathrm{C}$ is ill-conditioned and the error in the solution may become significant.
@@ V.1.2 Two-features model
where $r=R[{X^{(1)}},{X^{(2)}}]$ is the Pearson coefficient between the two features. When $r=0$, the two features are orthogonal, and $\mathrm{C}$ reduces to the identity matrix. When $r\to 1$ ($r\to-1$), the two features are perfectly correlated (anticorrelated), and hence ${\bf X}^{(1)}$ and ${\bf X}^{(2)}$ are linearly dependent.
@@ V.1.3 Origin of the oscillatory behavior of weights
We now seek to explain the oscillatory behavior of the weights observed in Fig. 3. This behavior arises from the ill-conditioning of the correlation matrix, because of the vanishing of its smallest eigenvalue $\lambda_{\rm min}$.
@@
V.2 Least squares regression revisited
@@
We can now build on the multicollinearity analysis to identify a range of optimal solutions within Ridge regression. To see how the regularization term in Eq. (32) mitigates the effect of multicollinearity, we compute the condition number in the Ridge setting. We get
@@
which avoids the divergence of $\kappa(\mathrm{C}+\alpha\mathrm{I})$ when multicollinearity is severe ($\lambda_{\rm min}\to 0$). In Fig. 6, we show $\kappa(\mathrm{C}+\alpha\mathrm{I})$ as a function of $\alpha$. In the range $\lambda_{\textrm{min}}\ll\alpha\ll\lambda_{\textrm{max}}$, the condition numbers decreases like $\lambda_{\textrm{max}}\alpha^{-1}$. To reduce the effects of multicollinearity to an acceptable degree, one should lower the condition numbers down to at least $10^{3}$ or less [48]. This occurs around $\alpha=0.1$. The regularization becomes meaningless for $\alpha\gtrsim\lambda_{\rm max}$, since it entirely suppresses correlations in $\mathrm{C}$. The crossover of $\kappa$ towards 1 occurs when $\alpha$ becomes of the order of $\lambda_{\textrm{max}}\approx 100$, as confirmed by Fig. 6. Since values of $\alpha$ in the range $0.1\lesssim\alpha\lesssim 100$ are acceptable but the performance accuracy decreases systematically in this range, see Fig. 5, we tentatively choose $\alpha=0.1$ as an optimal value. Finally, one can also see how the oscillatory behavior originating from multicollinearity is suppressed in Ridge regression. This point can already be grasped by comparing Fig. 3(a) and Fig. 3(b), but is best understood in the principal component basis, see the Appendix E.3.
@@ VI Towards interpretable linear regression models
The analysis presented in Sec. V shows that Ridge regression is an effective method to cope with the effects of multicollinearity in structure-dynamics datasets. One could then identify the most relevant features as those with the largest absolute weights, see Sec. VII. However, Ridge regression does not offer a principled way to perform feature selection: by definition, the model retains all the features. To achieve a simple physical picture, we must substantially reduce the dimensionality of the problem at hand. In this section, we consider a generalization of Ridge regression, called elastic net [72], that provides means to select the most relevant features, while reducing multicollinearity (Sec. VI.1). As an alternative approach, we use principal component regression, which builds upon a simple linear transformation of original features (Sec. VI.2).
VI.1 Elastic net and Lasso regression
@@ VI.1.1 Definition
Figure 7: Elastic net regression of the dynamic propensity using the BP descriptor. Main panel: Pearson coefficient $R[Y,\hat{Y}$] for the Lasso regression models (large circles) as a function of the number of selected features $P$. The values of $\alpha$ corresponding to the models with $P\leq 10$ are indicated in the figure. The small circles indicate results for values $(\alpha,\beta)$ with $\beta<1$. Inset: number of selected features $P$ as a function of $\beta$ for selected values of $\alpha$.
@@
$\alpha$ $P$ $R$ 0.53 0.53 0.53 0.51 0.47 0.39 0.35 0.31 0.28 -0.20 -0.28 -0.42
0.1 10 $0.74$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
0.2 8 $0.69$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
0.25 7 $0.64$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
0.3 5 $0.61$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
0.4 4 $0.60$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
0.45 3 $0.59$ $\bullet$ $\bullet$ $\bullet$
0.5 2 $0.57$ $\bullet$ $\bullet$
0.53 1 $0.53$ $\bullet$
@@
Table 2: Low-dimensional optimal models obtained using Lasso regression with the BP descriptor. The bullets indicate the most probable features selected by the models for selected values of $\alpha$. The numbers in italic below each feature are the Pearson coefficients between that feature and the dynamic propensity. The estimated statistical uncertainty on the Pearson coefficients is $\pm 0.01$.
VI.1.2 Model selection
To get a feeling for the role of $\beta$, we show in the inset of Fig. 7 the dependence of the number of selected features $P$ on $\beta$, for a few fixed values of $\alpha$. We average the results over five independent realizations of the train and test sets. As expected, the number of selected features decreases as the strength of the L1 regularization increases, reaching its minimum for $\beta=1$. A similar effect is observed with increasing $\alpha$. Note that a too large value of $\alpha$ may lead to the trivial solution $P=0$. For our dataset, we found that Lasso regression ($\beta=1$) requires $\alpha<0.53$ to yield a finite $P$.
In general, the prediction performance of the model will vary with $\alpha$ and $\beta$. The main panel of Fig. 7 shows the Pearson coefficient $R[Y,\hat{Y}$] between the solutions of several elastic net models and the dynamic propensity, as a function of $P$. The data for $\beta=1$ are indicated as large black symbols. We also include the results of a more extensive sampling of $(\alpha,\beta)$ pairs, shown as small dots. The envelope of $R$ vs. $P$ defines a set of optimal models, in the sense that they provide the best performance accuracy for a given size $P$ of the feature space. It is clear from the figure that these models coincide with those identified by Lasso regression. The typical values of the Pearson coefficients for the low-dimensional optimal models range from 0.53 for $P=1$, to 0.61 for $P=5$, to 0.74 for $P=10$. These values can be compared with the maximum value of 0.88 achieved when retaining the full descriptor, see also Sec. VII.
@@ VI.1.2 Model selection
In Table 2, we report the features selected by low-dimensional optimal models ($P\leq 10$) within Lasso regression. Note that these are the most probable selected features, estimated using 10 independent realizations of the train and test sets. We include the corresponding average Pearson coefficients with the dynamic propensity, both for the regression models and for the individual features. The low-dimensional models of Lasso regression successfully pinpoint the features that are most correlated to the dynamic propensity, found in either angular or radial sectors of the BP descriptor. While these models achieve a reasonable correlation with the dynamic propensity, they are still somewhat redundant. For instance, the features selected by the $P=2$ model, i.e., $\Psi^{SS}(20)$ and $G^{S}(11)$, are appreciably correlated with one another ($R=0.62$). Even worse, the models with $P=3,4,5$ and 7 contain two nearly identical angular features, $\Psi^{SS}(19)$ and $\Psi^{SS}(20)$, having mutual correlation $R=0.99$. This can be explained by the fact that for these models $\beta=1$, hence there is no suppression of multicollinearity. It may be possible to address this issue by using a minimum-redundancy-maximum-relevance selection scheme [74], which includes iteratively features that are highly correlated with the target, while penalizing those strongly correlated with the ones already taken. We leave an assessment of this approach to a future investigation.
@@ VI.2.1 Definition
$\frac{1}{N_{\mathcal{S}}}{\mathrm{X}^{\prime}}^{T}\mathrm{X}^{\prime}=\mathrm{U}^{T}\mathrm{C}\mathrm{U}=\mathrm{\Lambda},$ (45)
@@
Figure 8: Principal component analysis for the BP descriptor. (a) Eigenvalue $\lambda^{(f)}$ of each PC; (b) Pearson coefficient $R[\tilde{X}^{(f)},Y]$ between the dynamic propensity and each PC; (c) absolute value of the Pearson coefficient $|R[\tilde{X}^{(f)},Y]|$ between the dynamic propensity and each PC as a function of the eigenvalue $\lambda^{(f)}$ of the PC; (d) same as (c) but showing the eigenvalues in log scale. In panels (c) and (d), the top 5 PCs are marked by arrows.
@@ VI.2.2 Interpretation of the principal components
In Fig. 9 we report the Pearson coefficients between the projections on the first few PCs and the physically motivated features defined in Sec. III.4. Fluctuations along PC1 are moderately correlated to concomitant fluctuations of the local potential energy $\overline{u}$ and the bond orientational order parameter $\overline{\Psi}_{6}$. This structural mode gathers the bulk of the variance of the normalized dataset, but it is uncorrelated to the dynamic propensity. The interpretation of the second PC is sharp: structural fluctuations along PC2 are strongly correlated to the local number density $\overline{\rho}$ ($|R|=0.93$). Note that, as anticipated in Sec. III.6, in our model the fluctuations of the local packing fraction $\overline{\varphi}$ are strongly anti-correlated with those of $\overline{\rho}$. The remaining PCs do not seem to have clear connections with physically motivated features. Interestingly, none of the structural modes identified by PCA captures the structural fluctuations associated to $\overline{\Theta}$, which is most strongly correlated with the dynamic propensity. Note, however, a regression of $\overline{\Theta}$ against the full BP descriptor yields quite accurate results: the representation of this feature is simply spread over several features of the descriptor.
Figure 9: Pearson coefficients $R$ between the first five PCs and the physically motivated features. The latter are coarse-grained over a length $\ell=1.5$.